IMACS progress report no. 61
B. Bigelow 12 May 2003
I) Project Status
Currently scheduled design/fabrication tasks: Status:
a) Thermal testing, thermal enclosure Started
b) DSS testing (Birk/Hare) IP
c) Calibration sys. fab (Kowal) Done
d) CCD array testing (Thompson/Burley) IP
e) Dewar/Detector final integration (Thompson/Burley) IP
f) Grism procurement (Sutin) Done
g) Instrument electronics assembly (Bagish) Done
h) Control software design (Birk) on sched.
i) Short Camera opto-assembly (Bigelow) Done
j) IFU design and build (U. Durham) IP
Project events in April:
- IMACS PSR scheduled for 5/28/03 - external reviewers include:
Jerry Nelson, UCO/Lick
Dan Fabricant, Harvard/SAO
Robert Fata, Harvard/SAO
- Current voyage for shipment, Columbus Line #357, leaves Long Beach 7/5/03
- New center-field guider assembly nearly complete
- All 6 gratings now mounted in cells and ready for optical testing
- Last of three grisms (300 l/mm) delivered
Latest project images:
Final Assy: http://www.ociw.edu/instrumentation/imacs/devel/final/index.html
Disperser Server: http://www.ociw.edu/instrumentation/imacs/devel/dss/index.html
Filter Servers: http://www.ociw.edu/instrumentation/imacs/devel/fss/index.html
Thermal Testing http://www.ociw.edu/instrumentation/imacs/devel/th_testing/index.html
II) Optics (Sutin)
- Nathan Zhang Sutin, Born April 27, 2003, 6 pounds, 9 ounces
- Echellette prism due from Zygo 5/7/03
- 3rd fabrication effort for 300l/mm grism completed!
III) Electronics (Bagish)
1) RS-422 communications tested o.k. Spares ordered
2) 10 structure RTD thermal sensor cables rewired.
3) Rearranged sensor inputs to data acquisition unit and received surface mount
RTD sensors for cameras and collimator.
4) Shutter motors changed out to same type connectors for ease of future
5) Hard Stop lines engraved onto MCS Chassis 1-3 and face-plates reinstalled.
6) Hard limit LEDs added onto Chassis 4 and new faceplate installed.
7) Microphone cables audio lines installed.
8) Installed new SCDF motor driver (Gecko drive). Rebuilding focus cable with
properly keyed connector.
9) Obtained spare parts boxes.
1) Attach remaining RTD thermal sensors and pressure transducers.
2) Mount 2nd data acquisition unit in rack 2.
3) Backup DOS computer
4) Diagnose auto PC reboot.
5) Spare 80 pin cables
6) Spares list
7) Insulation of racks
8) Instruction manual
9) Book of documents
10) Diodes into chassis 4.
11) Check 5V at guide camera encoder
12) Insulate front and back of black ABS box.
13) Trouble-shoot reason for Chassis 3 triggering ground-loop protection circuitry.
14) Install shutter connector mounting brackets
15) Make up spare data cables.
16) Dress all cables
17) Attach instrument bulkhead shrouding
IV) Software (Birk/Clardy)
Testing all motion stages, cold testing starts 5/14/03
- Progress on astronomer GUI for mask cutting
- Optical models for telescope/instrument are being calibrated
- No mask cutting testing (waiting for LCO staff to make test masks).
V) Detector Systems (Thomson/Burley)
In April, Greg completed the redesign of the dsp and pci boards with
the revised optical link. The new design uses one fiber pair, and has
a more robust scrambler/De-scrambler in a bigger Altera chip. The
board designs have been sent out for manufacturing, and the bigger
Altera chips ordered.
Some progress on the image wings -- these are features that appear as
streaks on either side of a bright spot, and extend in the ccd
horizontal direction. Ian and Brian did a number of tests, including
rotating the dewar and imaging with different filters.
The wings show up with the spectroscopic clear filter, and appear the
same even when the dewar is rotated. The wings are gone when a V or R
filter is used. It seems that the wings are wavelength dependent
features, and are associated with the ccd (but not with the ccd
VI) Mechanics (Bigelow/Hare)
1) Continued plumbing installation
2) Re-designed SMS mounting plate for new CFG layout, fab completed
3) Re-assembled short camera with new S04 cell
4) Re-tested hydro-static flexure in Collimator doublet - redesigned C04 cell
5) Designed instrument air over-pressure system (dust protection)
6) Produced new slit mask frame design, first quote $3000
7) Shutter guards completed and mounted
8) Finished mounting all 6 gratings
1) IMACS pre-ship review 5/28/03! Produce review documentation
2) Complete design of air supply and exhaust plumbing
3) Re-assemble and re-install collimator with new C04 cell
4) Support mechanical and thermal testing
5) Explore new designs for slit mask frames for added stiffness
- Replaced damaged linear motor for f/2 camera shutter
- DSS wheel and drive testing under full load continuing
- New CFG design complete, drafting in progress, solids to Vince
- Minor re-design of SHG linear guides in progress
- New limit switches and mounting brackets in design for all guider stages
- Assembly of calibration lamp unit in progress
VII) Fabrication (Kowal)
1) Finished all Cable Tray parts.
2) Finished new Center Field Guider parts.
3) Final rework and modifications of parts as needed.
IX) Other Business
Next meeting: 10:15 AM, 6/4/03